OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line 3x + y = 7 and passes through the point (6, −1).

OpenStudy (jim766):

first need to find out the slope of 3x + y = 7 , get y by itself -3x -3x y = =3x + 7 Can you tell me this line's slope?

OpenStudy (jim766):

y = mx + b,,,,, the m is the slope

OpenStudy (jim766):

sorry typo, that is supposed to be y = -3x + 7

OpenStudy (anonymous):

Refer to the attached Mathematica result.